# Uncertainty-aware Distributional Offline Reinforcement Learning

### v1

@@ Abstract
Machine Learning, ICML
1 Introduction
In the domain of safety-critical applications, the practicality of implementing online RL is constrained due to the imperative for extensive random exploration, which may incur potential hazards or substantial costs. In response to this predicament, offline RL (Levine et al., 2020) endeavors to discern the most optimal policy based on pre-collected data. In safety-critical settings, there is a pronounced emphasis on cultivating uncertainty-averse decision-making while factoring in the inherent variability in performance. Regrettably, contemporary research in offline RL predominantly fixates on objectives associated with expected value, thereby neglecting the consideration of performance variability across distinct episodes.
There are two different types of uncertainties in offline RL: epistemic uncertainty and aleatoric uncertainty. Epistemic uncertainty is caused by the model itself and aleatoric uncertainty is caused by the environment (i.e., pre-collected data). Previous works address the epistemic uncertainty by using the risk-averse offline RL. The common approach is to use variational autoencoders (VAEs) to reconstruct the behavior policy, which can also help reduce bootstrapping errors  (Urpí et al., 2021; Lyu et al., 2022). However, these approaches heavily rely on imitation learning to align the learned policy with the empirical behavior policy, potentially leading to sub-optimal performance due to the presence of sub-optimal trajectories and the risk-neutral nature of the dataset (Ma et al., 2021). Offline RL leverages large datasets, but incorporating data from multiple behavioral policies presents challenges in maintaining fidelity to a high-quality policy (Ma et al., 2021; Wang et al., 2023; Kumar et al., 2022). Another alternative involves refining behavior policies (Ma et al., 2021), yet it lacks versatility for different environments or tasks. Additionally, the collected offline trajectories consist of a mixture of trajectories from various behavior policies. VAE-based methods struggle to distinguish these trajectories, resulting in sub-optimal performance. Furthermore, recent findings highlight VAE’s constraints in modeling complicated distributions (Pearce et al., 2023; Wang et al., 2023). In terms of the aleatoric uncertainty, it is caused by the environment stochasticity. It affects policy learning via the accumulated discounted returns. Previous works (Bellemare et al., 2017; Barth-Maron et al., 2018) propose to use the distributional offline RL to formulate the randomness where the distribution of the discounted cumulative reward is used instead of maximizing the expectation of accumulated discounted returns.
In this study, we address the challenge of handling both uncertainties simultaneously within the context of risk-averse offline RL. As discussed before, existing epistemic-aware methods (i.e., risk-averse offline RL) have the limitation of expressiveness caused by VAE. Inspired by the recent advancement of the diffusion model, which has a strong expressiveness when imitating human behaviors (Pearce et al., 2023), we design a novel off-policy RL algorithm method named Uncertainty-aware offline Distributional Actor-Critic (UDAC). UDAC leverages the controllable diffusion model as a key component in behavior policy modeling, offering several advantages: i). Elimination of Manual Behavior Policy Specification: UDAC eliminates the need for manually defined behavior policies, making it adaptable to a wide range of risk-sensitive RL tasks. ii). Accurate Modeling of the behavior policy: UDAC enhances the precision of modeling the full distribution of behavior policy through the controllable diffusion model, improving robustness in the face of environmental stochasticity (e.g., environment with unknown uncertainty). Moreover, in offline RL, data frequently derives from heterogeneous datasets, posing significant challenges in developing a robust policy model. For instance, datasets characterized by an overrepresentation of certain actions or containing atypically successful outcomes can lead to skewed perceptions of action efficacy. Such misconceptions may prompt the agent to overestimate the value of these actions, resulting in the formulation of suboptimal strategies. To address this issue, we employ the controllable diffusion model within UDAC to accurately represent the behavior policy. iii). Incorporation of Perturbation Model: UDAC goes beyond simple imitation learning by incorporating a perturbation model tailored to meet the requirements of risk-sensitive settings. The perturbation model aims to provide a certain level of uncertainty to empower the UDAC to effectively manage uncertainty.
Through extensive experiments on various benchmarks covering both risk-sensitive and risk-neutral scenarios: risk-sensitive D4RL, risky robot navigation, and risk-neutral D4RL. We demonstrate that UDAC outperforms most baselines in risk-sensitive D4RL and risky robot navigation. Moreover, UDAC achieves comparable performance to existing state-of-the-art methods in risk-neutral D4RL.
2 Related Work
@@
Pearce et al. (2023) suggest that the imitation of human behaviors can be improved through the use of expressive and stable diffusion models. Janner et al. (2022) employ a diffusion model as a trajectory generator in their Diffuser approach, where a full trajectory of state-action pairs forms a single sample for the diffusion model.  Singh et al. (2020) utilize a distributional critic in their algorithm, but it is restricted to the CVaR and limited to the online reinforcement learning (RL) setting. Furthermore, their use of a sample-based distributional critic renders the computation of the CVaR inefficient. To enhance the computational efficiency of the CVaR, Urpí et al. (2021) extend the algorithm to the offline setting. Ma et al. (2021) replace normal Q-learning with conservation offline Q-learning and an additional constant to penalize out-of-distribution (OOD) behaviors. However, Lyu et al. (2022) observe that the introduced extra constant is excessively severe and propose a more gentle penalized mechanism that uses a piecewise function to control when should apply the extra penalization.
@@
Differences between existing works. There are two similar works that also apply the diffusion model in offline reinforcement learning: Diffusion-QL (Wang et al., 2023) and Diffuser (Janner et al., 2022). Diffuser (Janner et al., 2022) is from the model-based trajectory-planning perspective, while Diffusion-QL (Wang et al., 2023) is from the offline model-free policy-optimization perspective. But our work is from the risk-sensitive perspective and extends it into the distributional offline reinforcement learning settings.
3 Background
@@
In this work, a Markov Decision Process (MDP) is considered with states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ that may be continuous. It includes a transition probability $P(\cdot|s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, reward function $R(\cdot|s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and a discount factor $\gamma\in[0,1)$. A policy $\pi(\cdot|s)$ is defined as a mapping from states to the distribution of actions. RL aims to learn $\pi(\cdot|s)$ such that the expected cumulative long-term rewards $\mathbb{E}_{s_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}[\sum_{t=0}^{\infty}\gamma^{t}R(\cdot|s_{t},a_{t})]$ are maximized. The state-action function $Q_{\pi}(s,a)$ measures the discounted return starting from state $s$ and action $a$ under policy $\pi$. We assume the reward function $R(\cdot|s,a)$ is bounded, i.e., $|R(\cdot|s,a)|\leq r_{max}$. Now, consider the policy $\pi(\cdot|s)$, the Bellman operator $\mathcal{T}^{\pi}:\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ is introduced by Riedmiller (2005) to update the corresponding $Q$ value,
@@
In distributional RL (Bellemare et al., 2017), the aim is to learn the distribution of discounted cumulative rewards, represented by the random variable $Z^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}R(\cdot|s_{t},a_{t})$. Similar to the $Q$ function and Bellman operator mentioned earlier, we can define a distributional Bellman operator as follows,
@@
4 Methodology
4.1 Overview of UDAC
In the critic network of UDAC, our aim is policy evaluation, and we rely on Equation 2 to achieve this objective. Following the approaches of Dabney et al. (2018a); Urpí et al. (2021); Ma et al. (2021), we implicitly represent the return through its quantile function, as demonstrated by Dabney et al. (2018a). In order to optimize the $Z$, we utilise a neural network with a learnable parameter $\theta$ to represent the quantile function. We express such implicit quantile function as $F_{Z}^{-1}(s,a;\tau)$ for $Z$, where $\tau\in[0,1]$ is the quantile level. Notice that, the original quantile function can only be used for discrete action. We can extend it into a continuous setting by considering all $s,a$, and $\tau$ as the inputs and only the quantile value as the output. To learn the $\theta$, the common strategy is the fitted distributional evaluation using a quantile Huber-loss $\mathcal{L}_{\mathcal{K}}$ Huber (1992). Our optimizing objective for the critic network is,
@@ 4.1 Overview of UDAC
where $(s,a,r,s^{\prime})\sim\mathcal{D}$ and $a^{\prime}\sim\pi(\cdot|s^{\prime})$. The $\delta$ is also known as the distributional TD error (Dabney et al., 2018b). And the quantile Huber-loss $\mathcal{L}_{\mathcal{K}}$ is represented as,
@@
where $\mathcal{K}$ is the threshold. Notice that there is a hyper-parameter $\tau$ which may affect our objective function. Here, we use a simple strategy to control the $\tau$: Uniformly random sampling $\tau\sim U(0,1)$. It is worth mentioning that there is an alternative way for controlling $\tau$, which is a quantile level proposal network $P_{\psi}(s,a)$ Yang et al. (2019). To simplify the overall model and reduce the number of hyperparameters, we just use the simplified random sampling approach and leave the quantile proposal network for future work. The overall loss function for the critic is
@@
where $N,K$ are the numbers of quantile levels. In the risk-sensitive setting, deterministic policies are favoured over stochastic policies due to the latter’s potential to introduce additional randomness, as noted by Pratt (1978). Moreover, in offline settings, there is no benefit to exploration commonly associated with stochastic policies. Therefore, the focus is on parameterized deterministic policies, represented by $\pi_{\phi}(s):\mathcal{S}\rightarrow\mathcal{A}$. Given the risk-distortion function $\mathbb{D}$, the loss function of actor is written as,
@@
with the marginal state distribution $\rho_{b}$ from the behavior policy. To optimize it, we can backpropagate it through the learned critic at the state that sampled from the behavior policy. It is equivalent to maximising risk-averse performance. Then, combining the above process with the SAC (Haarnoja et al., 2018) leads to our novel UDAC model.
4.2 Generalized Behavior Policy Modeling
We separate the actor into two components: an imitation component (Fujimoto et al., 2019) and a conditionally deterministic perturbation model $\xi$,
@@ 4.2 Generalized Behavior Policy Modeling
where $\xi(\cdot|s,\beta)$ can be optimized by using the actor loss function(i.e., $-\mathbb{E}_{s\sim\rho_{b}}[\mathbb{D}(F_{Z}^{-1}(s,a;\tau))]$) that was defined previously, and $\beta$ is the action that sampled from behavior policy $\pi_{b}$, $\lambda$ is a scale that used to control the magnitude of the perturbation. As aforementioned, we prefer the deterministic policy over the stochastic policy because of the randomness. However, the bootstrapping error will appear on the offline RL setting. Existing works (Urpí et al., 2021; Wu et al., 2019; Ma et al., 2021; Lyu et al., 2022) use policy regularization to regularize how far the policy can deviate from the behavior policy. But the policy regularization will adversely affect the policy improvement by limiting the exploration space, and only suboptimal actions will be observed. Moreover, the fixed amount of data from the offline dataset may not be generated by a single, high-quality behaviour policy and thus affect the imitation learning performance (i.e., the logged data may come from different behavior policies). Hence, behavioral cloning (Pomerleau, 1991) would not be the best choice. Moreover, behavioral cloning suffers from mode-collapse, which will affect actor’s optimization. Here, we attempt to use the generative method to model the behaviour policy from the observed dataset. The most popular direction is the cVAE-based approaches (Urpí et al., 2021; Wu et al., 2019) which leverages the capability of the auto-encoder to reconstruct the behavior policy. However, these methods did not account for the noisy behavior policies resulting from the offline reinforcement learning environment. This environment typically comprises datasets collected from various agents, each potentially exhibiting distinct modalities. This suggests that certain behavior policies within these datasets might be sub-optimal, adversely affecting overall performance. To address this, we seek denoising diffusion probabilistic model (DDPM) (Ho et al., 2020) to facilitate robust behavior policy modelling. The diffusion model facilitates a gradual lossy decompression approach, which extends autoregressive decoding. Thus, UDAC utilizes this model as a behavior policy to utilize pre-gathered sub-optimal trajectories for final policy learning.
@@
Following the approach of Ho et al. (2020), we set $\epsilon$ to 0 when $i=1$ to improve the sampling quality. The objective function used to train the conditional $\epsilon$-model is,
$\mathcal{L}_{d}(\psi)=$
@@
where $\mathcal{U}$ is a uniform distribution over the discrete set, denoted as $\{1,\cdots,N\}$. This diffusion model loss $\mathcal{L}_{d}(\psi)$ aims to learn from the behavior policy $\pi_{b}$ instead of cloning it. $\mathcal{L}_{d}(\psi)$ can be optimized by sampling from single diffusion step $i$ for each data point. However, it will be affected significantly by $N$, which is known as the bottleneck of the diffusion (Kong & Ping, 2021). We add this term to the actor loss and construct the final loss function. We then replace the term $\pi_{b}$ in the Equation 7 with $d_{\psi}(s,a)$ (i.e., the diffusion behavior policy representation).
4.3 Controllable Diffusion Process
@@
We can further simplify $p(s|a^{i-1},a^{i})=p(s|a^{i-1})$ via conditional independence assumptions from prior work on controlling diffusions (Song et al., 2020). Consequently, for the $i$-step, we run gradient update on $a^{i-1}$,

### v2

# Diffusion Policies for Risk-Averse Behavior Modeling in Offline Reinforcement Learning

@@ Abstract
I INTRODUCTION
In safety-critical applications, deploying online reinforcement learning (RL) is often infeasible due to the extensive exploration required, which can result in hazardous or costly outcomes. Offline RL [1] addresses this issue by enabling the learning of optimal policies from pre-collected data, thereby mitigating the risks associated with direct exploration. However, safety-critical environments require policies that not only perform well but also manage risk effectively. Risk in these settings arises from the unpredictability of the environment’s behavior or the generalization ability of the learned policy in unseen situations, making risk-aware decision-making essential.
In most existing offline RL approaches, the primary focus is on maximizing expected returns [1], often without adequately accounting for the variability and potential adverse consequences of certain actions. This risk-neutral perspective can result in suboptimal performance, particularly in scenarios where avoiding worst-case outcomes is critical. Uncertainty in the quality and consistency of the collected data in offline RL is closely associated with risk, as unreliable or inconsistent data can lead to the development of policies that perform poorly in practice.
To address these risks, current risk-aware methods, such as those based on conditional variational autoencoders (cVAEs), attempt to mitigate bootstrapping errors by reconstructing behavior policies [2, 3]. However, these methods frequently rely on imitation learning, which aligns the learned policy with the observed behavior in the dataset. This dependence makes them susceptible to suboptimal performance when the dataset contains diverse or suboptimal trajectories [4]. Furthermore, cVAEs are limited in their ability to model complex behavior distributions, a limitation that becomes particularly problematic in offline RL settings where heterogeneous datasets—combining trajectories from multiple behavior policies—are common. These heterogeneous datasets present significant challenges in developing policies that are both high-quality and risk-aware [5, 6]. While one such method attempts to refine behavior policies [4], it often lacks the flexibility required to generalize effectively across different environments and tasks, further complicating risk management.
To address the aforementioned challenges, it is essential to account for uncertainty when managing risk, as it reflects both the unpredictability of the environment and the quality of the pre-collected data. In response, we propose Uncertainty-aware offline Distributional Actor-Critic (UDAC), a novel approach that integrates uncertainty-aware policy learning with risk management. UDAC leverages diffusion models, which have demonstrated strong capabilities in modeling behavior policies [5], to address both risk and uncertainty within a unified framework.
Our method introduces several key innovations: (i) Elimination of Manual Behavior Policy Specification: UDAC automates the modeling of behavior policies, removing the need for manual design and making it adaptable to a wide variety of risk-sensitive tasks. (ii) Accurate Behavior Policy Modeling: By utilizing a controllable diffusion model, UDAC accurately captures the full distribution of behavior policies, which is critical for robust decision-making in environments with substantial uncertainty and risk. iii) UDAC incorporates a perturbation model tailored for risk-sensitive settings, allowing it to balance exploration and caution. This model introduces controlled variability into the policy, enabling UDAC to manage risk more effectively by accounting for uncertainties in both the environment and the dataset.
Extensive experimental evaluations on benchmarks, including risk-sensitive D4RL, risky robot navigation, and risk-neutral D4RL, demonstrate that UDAC surpasses baseline methods in risk-sensitive tasks while maintaining competitive performance in risk-neutral settings. These results indicate that UDAC offers a comprehensive framework for managing both risk and uncertainty in offline RL, positioning it as an effective tool for navigating complex, safety-critical environments.
II Related Work
@@
[5] suggest that the imitation of human behaviors can be improved through the use of expressive and stable diffusion models. [7] employ a diffusion model as a trajectory generator in their Diffuser approach, where a full trajectory of state-action pairs forms a single sample for the diffusion model.  [8] utilize a distributional critic in their algorithm, but it is restricted to the CVaR and limited to the online reinforcement learning (RL) setting. Furthermore, their use of a sample-based distributional critic renders the computation of the CVaR inefficient. To enhance the computational efficiency of the CVaR, [2] extend the algorithm to the offline setting. [4] replace normal Q-learning with conservation offline Q-learning and an additional constant to penalize out-of-distribution (OOD) behaviors. However, [3] observe that the introduced extra constant is excessively severe and propose a more gentle penalized mechanism that uses a piecewise function to control when applying the extra penalization. [9] focuses on the safe offline RL domain which also uses the risk-averse approach.
@@
Differences between existing works. There are two similar works that also apply the diffusion model in offline reinforcement learning: Diffusion-QL [6] and Diffuser [7]. Diffuser [7] is from the model-based trajectory-planning perspective, while Diffusion-QL [6] is from the offline model-free policy-optimization perspective. However, our work is from the risk-sensitive perspective and extends it into the distributional offline reinforcement learning settings.
III Background
@@
In this work, a Markov Decision Process (MDP) is considered with states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ that may be continuous. It includes a transition probability $P(\cdot|s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$, reward function $R(\cdot|s,a):\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and a discount factor $\gamma\in[0,1)$. A policy $\pi(\cdot|s)$ is defined as a mapping from states to the distribution of actions. RL aims to learn $\pi(\cdot|s)$ such that the expected cumulative long-term rewards $\mathbb{E}_{s_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}[\sum_{t=0}^{\infty}\gamma^{t}R(\cdot|s_{t},a_{t})]$ are maximized. The state-action function $Q_{\pi}(s,a)$ measures the discounted return starting from state $s$ and action $a$ under policy $\pi$. We assume the reward function $R(\cdot|s,a)$ is bounded, i.e., $|R(\cdot|s,a)|\leq r_{max}$. Now, consider the policy $\pi(\cdot|s)$, the Bellman operator $\mathcal{T}^{\pi}:\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}\rightarrow\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ is introduced by [10] to update the corresponding $Q$ value,
@@
In distributional RL [11], the aim is to learn the distribution of discounted cumulative rewards, represented by the random variable $Z^{\pi}(s,a)=\sum_{t=0}^{\infty}\gamma^{t}R(\cdot|s_{t},a_{t})$. Similar to the $Q$ function and Bellman operator mentioned earlier, we can define a distributional Bellman operator as follows,
@@
IV Methodology
IV-A Overview of UDAC
In the critic network of UDAC, our aim is policy evaluation, and we rely on Equation 2 to achieve this objective. Following the approaches of [12, 2, 4], we implicitly represent the return through its quantile function, as demonstrated by [12]. In order to optimize the $Z$, we utilise a neural network with a learnable parameter $\theta$ to represent the quantile function. We express such implicit quantile function as $F_{Z}^{-1}(s,a;\tau)$ for $Z$, where $\tau\in[0,1]$ is the quantile level. Notice that, the original quantile function can only be used for discrete action. We can extend it into a continuous setting by considering all $s,a$, and $\tau$ as the inputs and only the quantile value as the output. To learn the $\theta$, the common strategy is the fitted distributional evaluation using a quantile Huber-loss $\mathcal{L}_{\mathcal{K}}$ [13]. Our optimizing objective for the critic network is,
@@ IV-A Overview of UDAC
where $(s,a,r,s^{\prime})\sim\mathcal{D}$ and $a^{\prime}\sim\pi(\cdot|s^{\prime})$. The $\delta$ is also known as the distributional TD error [14]. And the quantile Huber-loss $\mathcal{L}_{\mathcal{K}}$ is represented as,
@@
where $\mathcal{K}$ is the threshold. Notice that there is a hyper-parameter $\tau$ which may affect our objective function. Here, we use a simple strategy to control the $\tau$: Uniformly random sampling $\tau\sim U(0,1)$. It is worth mentioning that there is an alternative way for controlling $\tau$, which is a quantile level proposal network $P_{\psi}(s,a)$ [15]. To simplify the overall model and reduce the number of hyperparameters, we just use the simplified random sampling approach and leave the quantile proposal network for future work. The overall loss function for the critic is
@@
where $N,K$ are the numbers of quantile levels. In the risk-sensitive setting, deterministic policies are favoured over stochastic policies due to the latter’s potential to introduce additional randomness, as noted by [16]. Moreover, in offline settings, there is no benefit to exploration commonly associated with stochastic policies. Therefore, the focus is on parameterized deterministic policies, represented by $\pi_{\phi}(s):\mathcal{S}\rightarrow\mathcal{A}$. Given the risk-distortion function $\mathbb{D}$, the loss function of actor is written as,
@@
with the marginal state distribution $\rho_{b}$ from the behavior policy. To optimize it, we can backpropagate it through the learned critic at the state that sampled from the behavior policy. It is equivalent to maximising risk-averse performance. Then, combining the above process with the SAC [17] leads to our novel UDAC model.
IV-B Generalized Behavior Policy Modeling
We separate the actor into two components: an imitation component [18] and a conditionally deterministic perturbation model $\xi$,
@@ IV-B Generalized Behavior Policy Modeling
where $\xi(\cdot|s,\beta)$ can be optimized by using the actor loss function(i.e., $-\mathbb{E}_{s\sim\rho_{b}}[\mathbb{D}(F_{Z}^{-1}(s,a;\tau))]$) that was defined previously, and $\beta$ is the action that sampled from behavior policy $\pi_{b}$, $\lambda$ is a scale that used to control the magnitude of the perturbation. As aforementioned, we prefer the deterministic policy over the stochastic policy because of the randomness. However, the bootstrapping error will appear on the offline RL setting. Existing works [2, 19, 4, 3] use policy regularization to regularize how far the policy can deviate from the behavior policy. But the policy regularization will adversely affect the policy improvement by limiting the exploration space, and only suboptimal actions will be observed. Moreover, the fixed amount of data from the offline dataset may not be generated by a single, high-quality behaviour policy and thus affect the imitation learning performance (i.e., the logged data may come from different behavior policies). Hence, behavioral cloning [20] would not be the best choice. Moreover, behavioral cloning suffers from mode-collapse, which will affect actor’s optimization. Here, we attempt to use the generative method to model the behaviour policy from the observed dataset. The most popular direction is the cVAE-based approaches [2, 19] which leverages the capability of the auto-encoder to reconstruct the behavior policy. However, these methods did not account for the noisy behavior policies resulting from the offline reinforcement learning environment. This environment typically comprises datasets collected from various agents, each potentially exhibiting distinct modalities. This suggests that certain behavior policies within these datasets might be sub-optimal, adversely affecting overall performance. To address this, we seek denoising diffusion probabilistic model (DDPM) [21] to facilitate robust behavior policy modelling. The diffusion model facilitates a gradual lossy decompression approach, which extends autoregressive decoding. Thus, UDAC utilizes this model as a behavior policy to utilize pre-gathered sub-optimal trajectories for final policy learning.
@@
Following the approach of [21], we set $\epsilon$ to 0 when $i=1$ to improve the sampling quality. The objective function $\mathcal{L}_{d}(\psi)$ used to train the conditional $\epsilon$-model is,
@@
where $\mathcal{U}$ is a uniform distribution over the discrete set, denoted as $\{1,\cdots,N\}$. This diffusion model loss $\mathcal{L}_{d}(\psi)$ aims to learn from the behavior policy $\pi_{b}$ instead of cloning it. $\mathcal{L}_{d}(\psi)$ can be optimized by sampling from single diffusion step $i$ for each data point. However, it will be affected significantly by $N$, which is known as the bottleneck of the diffusion [22]. We add this term to the actor loss and construct the final loss function. We then replace the term $\pi_{b}$ in the Equation 7 with $d_{\psi}(s,a)$ (i.e., the diffusion behavior policy representation).
@@
We can further simplify $p(s|a^{i-1},a^{i})=p(s|a^{i-1})$ via conditional independence assumptions from prior work on controlling diffusions [23]. Consequently, for the $i$-step, we run gradient update on $a^{i-1}$,